# Stability-Aware Training of Neural Network Interatomic Potentials with Differentiable Boltzmann Estimators

### v1

@@ Abstract
Neural network interatomic potentials (NNIPs) are an attractive alternative to ab-initio methods for molecular dynamics (MD) simulations. However, they can produce unstable simulations which sample unphysical states, limiting their usefulness for modeling phenomena occurring over longer timescales. To address these challenges, we present Stability-Aware Boltzmann Estimator (StABlE) Training, a multi-modal training procedure which combines conventional supervised training from quantum-mechanical energies and forces with reference system observables, to produce stable and accurate NNIPs. StABlE Training iteratively runs MD simulations to seek out unstable regions, and corrects the instabilities via supervision with a reference observable. The training procedure is enabled by the Boltzmann Estimator, which allows efficient computation of gradients required to train neural networks to system observables, and can detect both global and local instabilities. We demonstrate our methodology across organic molecules, tetrapeptides, and condensed phase systems, along with using three modern NNIP architectures. In all three cases, StABlE-trained models achieve significant improvements in simulation stability and recovery of structural and dynamic observables. In some cases, StABlE-trained models outperform conventional models trained on datasets 50 times larger. As a general framework applicable across NNIP architectures and systems, StABlE Training is a powerful tool for training stable and accurate NNIPs, particularly in the absence of large reference datasets.
Molecular dynamics (MD) simulation is a staple method of computational science, enabling high-resolution spatiotemporal modeling of atomistic systems throughout biology, chemistry, and materials science [undeft]. Under the Born-Oppenheimer approximation, system evolution is governed by the underlying potential energy surface (PES), which is a function of the nuclear Cartesian coordinates [undefj]. While the atomic forces needed for MD simulation can be obtained on-the-fly via ab-initio quantum-mechanical (QM) calculations [undefk], the unfavorable scaling of this approach makes it prohibitively expensive for realistic system sizes and timescales [undefu].
There is a long history of using machine learning (ML) approaches in place of ab-initio methods to efficiently approximate the global PES [undeff, undefe, undefa, undefaab]. Neural network interatomic potentials (NNIP) have recently emerged as a highly successful class of ML methods, demonstrating the ability to capture complex many-body interactions, and in some cases transfer flexibly across chemical space [undefaae, undefab, undefao, undefw, undefx, undefd, undefat, undefc, undefaaf]. NNIPs, typically parameterized as graph neural networks [undefaac, undefaf], are trained by matching energy and forces of a molecule or material from a reference dataset of QM calculations, such as Density Functional Theory (DFT) [undefad]. NNIPs trained on large ab-initio datasets are increasingly being used to model challenging and important chemical systems with favorable results [undefar, undefaj, undefn, undefaad, undefaak, undefap, undefm, undefb, undefai, undefaag, undefy, undefr].
MD simulations aim to accurately estimate system observables like the radial distribution function, virial stress tensor, and diffusivity coefficient. This often requires long simulation to fully explore the underlying PES and minimize sampling error. Unfortunately, NNIPs are known to produce unstable simulations, meaning that they can irreversibly enter unphysical regions of phase space (e.g., a bond breaking event in a non-reactive system at low temperature) [undefv, undefaaj, undefaap, undefg, undefas, undefaas]. Sampling of such regions often quickly leads to unrecoverable simulation collapse as the NNIP drifts far from the distribution of its training data. As a result, computed observables may have high statistical error due to insufficient sampling, or be altogether inaccurate. This can limit the ability of NNIP-based MD simulations to investigate long-timescale phenomena like ion diffusion and protein folding, as well as rare events that may require extensive sampling to encounter. Figure 1 illustrates typical NNIP instability behaviors and provides examples of unstable structures sampled during MD simulation.
Insufficient coverage of phase space and inaccuracies within the reference datasets utilized for training NNIPs have been suggested as factors contributing to instability [undefv, undefaaj, undefaap, undefg, undefas, undefaas]. Expanding the dataset requires additional QM calculations, and improving its accuracy requires more exact QM approaches [undefal]. Both options are very computationally expensive, particularly for larger systems like macromolecules. Recent work has even suggested that expanding the dataset may not always improve stability [undefv], consistent with the notion that phase space is sampled with exponentially vanishing probability based on the local potential energy. This suggests the need for additional sources of information beyond energies and forces to train stable NNIPs.
@@
Figure 1: Neural network interatomic potential (NNIP) failure modes. (a) Illustrative examples of true and learned potential energy surfaces (PES) and resulting dynamics for unstable NNIPs. NNIPs can be highly accurate in approximating most of the PES, but contain localized ”holes” where energy and forces estimates deviate significantly from the true PES, leading to sampling of highly unphysical regimes. Once such a regime is reached, further simulation typically does not re-enter physically plausible regions. As a result, observables computed from MD simulation may be inaccurate or have high statistical error. (b) Examples of stable versus unstable configurations sampled by NNIPs during molecular dynamics simulation of systems considered in this work. The unstable configurations contain unphysical bond lengths, angles, and coordination structures which are not meaningful for observable calculations. (c) Selected states from an unstable MD trajectory of aspirin. A unidirectional progression is demonstrated, in which a stable, physical configuration evolves into a highly unstable, unphysical configuration.
We posit that system observables derived from experiment or high-fidelity simulation can serve this purpose. Parameterizing NNIPs to match observables is an alternative approach to conventional energy-and-forces-training [undefaal, undefaaq]. This approach has been successful in learning potentials of condensed-phase [undefaar] and titanium [undefaaa] systems, enhanced sampling of rare events [undefaah], and protein folding simulation [undefac, undefau, undefag]. However, using observables to improve the stability of NNIPs in downstream MD simulations remains unexplored.
@@
In this work, we introduce Stability-Aware Boltzmann Estimator (StABlE) Training, a procedure designed to produce NNIPs that are both stable and accurate. The core idea behind StABlE is to use MD simulation to rapidly explore regions of molecular phase space where the NNIP becomes unstable, followed by a targeted refinement of these regions using reference system observables. The key to efficient and numerically stable training lies in the Boltzmann Estimator, which avoids computing gradients directly through MD simulations. StABlE Training is a self-contained, efficient process that leverages learning signals from both reference observables and existing QM data, with no reliance on performing additional QM reference calculations. To our knowledge, StABlE Training is the first method that combines these data modalities to improve the stability of neural network potentials in MD simulations. Its flexibility allows for integration with any NNIP architecture or molecular system, making it a valuable addition to existing NNIP workflows.
@@
We demonstrate StABlE Training on three systems and NNIP architectures: 1) simulation of aspirin with SchNet [undefaae], 2) simulation of the Ac-Ala3-NHMe tetrapeptide with NequIP [undefd], and 3) simulation of an all-atom water system with GemNet-T [undefw]. Relative to NNIPs trained solely on reference energy and forces data, our StABlE-trained models produce significantly more stable downstream MD simulations and recover both structural and dynamical observables more accurately. For an aspirin simulation with a SchNet model trained on 1,000 reference structures, StABlE Training increases the median stability from 35 to 140 picoseconds, exhibits better generalization to unseen simulation temperatures, and outperforms models trained on 50 times larger, labeled datasets. For an all-atom water simulation with a GemNet-T model trained on 1,000 reference structures, StABlE Training improves the median stability from 23 to 52 picoseconds and yields an 87% accuracy improvement in estimation of the diffusivity coefficient. Our results suggest that utilizing both quantum-mechanical and observable-based modalities is required to fully exploit the available learning signal in reference datasets and train stable and accurate NNIPs.
@@ 1 Preliminaries
Molecular dynamics simulation is used to evolve the positions and momenta of an atomistic system. Consider a system of $N$ atoms. The state of the system at time $t$ is defined by $\Gamma(t)=\{r(t),p(t)\}$, where $r(t),p(t)\in\mathbb{R}^{N\times 3}$ are the position and momenta of the atoms. We assume the system has a scalar-valued Hamiltonian function $\mathcal{H}:\mathbb{R}^{N\times 3}\times\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}$ of the form,
$\mathcal{H}(\Gamma)=\sum_{i=1}^{N}\frac{p_{i}^{2}}{2m_{i}}+U(r)\,$, (1)
where $U:\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}$ is a potential energy function and $m_{i}$ and $p_{i}$ are the mass and momentum of atom $i$. The system obeys Hamilton’s laws of motion,
$\dot{r_{i}}=\frac{\partial H}{\partial p_{i}}=\frac{p_{i}}{m_{i}}\,,\quad\dot{p_{i}}=-\frac{\partial H}{\partial r_{i}}=-\frac{\partial U(r_{i})}{\partial r_{i}}\,.$ (2)
Evolving these equations in time for $T_{f}$ steps produces a sequence of simulation states, $\bar{\Gamma}=\{\Gamma(t)\}_{t=0}^{T_{f}}$, which conserve the Hamiltonian and constitute samples from the microcanonical ensemble. It is also common to couple a thermostat to the simulation to conserve temperature instead of the Hamiltonian, thereby sampling from the canonical ensemble. In this case, the state distribution has a Boltzmann form: $P(\Gamma)\propto e^{-\frac{H(\Gamma)}{k_{B}T}}$ where $T$ is the temperature and $k_{B}$ is Boltzmann’s constant.
Training Neural Network Potentials with Reference Energies and Forces.
In the context of neural network interatomic potentials (NNIPs), the potential energy is approximated by a function $U_{\theta}$ with learnable parameters $\theta$ given by a neural network. When training a model to learn from quantum mechanical reference data, $\theta$ is obtained by using stochastic gradient descent (SGD) to minimize the following energy and forces loss function:
$\mathcal{L}_{\text{QM}}(\theta)\stackrel{{\scriptstyle\text{def}}}{{=}}$ $\mathbb{E}_{\Gamma\sim P_{\text{QM}}(\Gamma)}\Big[\lambda_{U}\|U_{\text{ref}}(\Gamma)-U_{\theta}(\Gamma)\|_{2}^{2}$ (3)
@@
$\quad+\lambda_{F}||F_{\text{ref}}(\Gamma)+\nabla_{r}{U_{\theta}(\Gamma)}||_{2}^{2}\Big]\,$,
where $P_{\text{QM}}(\Gamma)$ is the equilibrium distribution induced by ab-initio (or high-fidelity) simulation, $U_{\text{ref}}$ and $F_{\text{ref}}$ are the high-fidelity reference energy and force functions, and $\lambda_{U}$ and $\lambda_{F}$ are hyperparameters controlling the relative weights of the energy and force loss terms. Note that while we have adopted the notation $U(\Gamma)$ for convenience, the potential energy is a function only of the position component, $r$, of the state $\Gamma$. In this work, $P_{\text{QM}}(\Gamma)$ is approximated by a reference dataset $\mathcal{D}_{\text{train}}$ of states sampled from a high-fidelity simulation, such as ab-initio molecular dynamics. We note that in some cases, $\mathcal{D}_{\text{train}}$ may not be fully representative of $P_{\text{QM}}(\Gamma)$, due to biased sampling procedures favoring rare or interesting events.
Estimation of System Observables from MD Simulation.
@@
Observables, $g(\Gamma(t))$, characterize the state of a MD simulation at time $t$, and relate to macroscopic properties or experimental measurements of the system. Examples include the radial and angular distribution functions, velocity autocorrelation function, and diffusivity coefficient. Observables can be computed conveniently as an empirical average over states from a MD simulation. This is justified by the ergodicity hypothesis, under which a time-average over a sufficiently long simulation converges to a distributional average over the Boltzmann distribution. For more detail on the observables used in this work, please refer to Supplementary Section A.4.
@@ 2 Methods
We present StABlE Training, our proposed procedure to train stable and accurate NNIPs by leveraging both reference quantum mechanical data and system observables. Our approach is enabled by Boltzmann Estimators, which allows gradient-based optimization of NNIPs based on system observables.
@@ 2.1 Boltzmann Estimator
Reference system observables can be estimated from high-fidelity MD simulations or experimental measurements. Since observables are linked to the NNIP parameters through a molecular simulation, training with this source of information requires a reliable way to optimize through MD trajectories.
Figure 2: Schematic of Stability-Aware Boltzmann Estimator (StABlE) Training procedure. Our proposed StABlE Training procedure begins with conventional pre-training on a small reference dataset of QM calculations (in the loss, $\hat{U}$ refers to the NNIP-predicted energy). This dataset remains fixed throughout the procedure, and is never expanded with new calculations. Upon convergence of pre-training, StABlE then alternates between two main phases, simulation and learning. In the simulation phase, we perform many molecular dynamics simulations in parallel with the NNIP and find regions of instability. When a sufficient fraction of simulations become unstable, we enter the learning phase, where the NNIP is further trained to match known system observables from ab-initio simulation or experiment (in the loss, $\hat{g}$ refers to the NNIP-predicted observable). After a sufficient reduction in the portion of unstable trajectories, we re-enter the simulation phase, and repeat the training cycle until a predetermined computational budget is reached. Sample stable and unstable molecules are included next to the corresponding graphs.
Formally, we define $g(\Gamma)$ to be a vector-valued observable of a state $\Gamma$, and $g_{\text{ref}}$ to be the reference value of the observable. To train our NNIP with information about the system observable, we seek to minimize the following loss function:
@@
where $P_{\theta}(\Gamma)$ is the equilibrium distribution induced by the NNIP $U_{\theta}$. In this work, we consider systems with a fixed volume, temperature, and number of particles, corresponding to the canonical ensemble. This means that microstates obey a Boltzmann distribution, $P_{\theta}(\Gamma)={\exp\left(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma)\right)}/{C(\theta)}$, where the Hamiltonian $\mathcal{H}_{\theta}(\Gamma)=\sum_{i=1}^{N}\frac{p_{i}^{2}}{2m_{i}}+U_{\theta}(\Gamma)$, $T$ is the sampling temperature, $k_{B}$ is Boltzmann’s constant, and $C(\theta)=\int\exp(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma^{\prime}))\mathop{}\!\mathrm{d}\Gamma^{\prime}$ is the normalizing partition function. MD simulation is required at each training iteration to sample from this distribution and estimate the loss.
We optimize the parameters of the NNIP, $\theta$, using stochastic gradient descent. This requires computing gradients of Equation 4 with respect to $\theta$. The required gradient, $\nabla_{\theta}\mathcal{L}_{\text{obs}}$, can be decomposed via the chain rule as follows:
$\nabla_{\theta}\mathcal{L}_{\text{obs}}^{\top}$ $=\frac{\partial\mathcal{L}_{\text{obs}}}{\partial{\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)[g(\Gamma)]}}}\frac{\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}{\partial\theta}$ (5)
The non-trivial quantity to compute is the Jacobian, ${\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}/{\partial\theta}$. One way to estimate this quantity is by using a chain rule expansion that corresponds to each step of the unrolled MD simulation [undefaaq]. However, this approach is memory-intensive and can lead to numerical instability for long trajectories [undefaar, undefaah]. We can instead leverage the known Boltzmann form of the state distribution to construct an unbiased estimator of the Jacobian.
@@
Given $N$ independent samples $\Gamma^{(1)},\ldots,\Gamma^{(N)}$ from a Boltzmann distribution $P_{\theta}(\Gamma)$, we define the $N$-sample Boltzmann estimator $\mathcal{E}(\Gamma^{(1)},\ldots,\Gamma^{(N)})$ as,
$\mathcal{E}(\Gamma^{(1)},\ldots,\Gamma^{(N)})=$ (6)
$\frac{N}{k_{B}T(N-1)}\left[\hat{\mathbb{E}}\left[g(\Gamma)\right]\hat{\mathbb{E}}\big[{\nabla_{\theta}U_{\theta}(\Gamma)}\right]^{\top}\!-\hat{\mathbb{E}}\left[{g(\Gamma)\cdot\nabla_{\theta}U_{\theta}(\Gamma)}^{\top}\right]\big]\,$,
where $\hat{\mathbb{E}}[f(\Gamma)]=\frac{1}{N}\sum_{i=1}^{N}f(\Gamma^{(i)})$ denotes an empirical mean over the samples.
This estimator provides an unbiased estimate of the Jacobian $\frac{\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}{\partial\theta}$. Our proof is provided in Supplementary Section A.1. This result can also be derived using thermodynamic perturbation theory [undefaau, undefaal].

### v2

# mddefDefinition

@@ Abstract
Machine learning force fields (MLFFs) are an attractive alternative to ab-initio methods for molecular dynamics (MD) simulations. However, they can produce unstable simulations, limiting their ability to model phenomena occurring over longer timescales and compromising the quality of estimated observables. To address these challenges, we present Stability-Aware Boltzmann Estimator (StABlE) Training, a multi-modal training procedure which leverages joint supervision from reference quantum-mechanical calculations and system observables. StABlE Training iteratively runs many MD simulations in parallel to seek out unstable regions, and corrects the instabilities via supervision with a reference observable. We achieve efficient end-to-end automatic differentiation through MD simulations using our Boltzmann Estimator, a generalization of implicit differentiation techniques to a broader class of stochastic algorithms. Unlike existing techniques based on active learning, our approach requires no additional ab-initio energy and forces calculations to correct instabilities. We demonstrate our methodology across organic molecules, tetrapeptides, and condensed phase systems, using three modern MLFF architectures. StABlE-trained models achieve significant improvements in simulation stability, data efficiency, and agreement with reference observables. By incorporating observables into the training process alongside first-principles calculations, StABlE Training can be viewed as a general semi-empirical framework applicable across MLFF architectures and systems. This makes it a powerful tool for training stable and accurate MLFFs, particularly in the absence of large reference datasets.
1 Introduction
Molecular dynamics (MD) simulation is a staple method of computational science, enabling high-resolution spatiotemporal modeling of atomistic systems throughout biology, chemistry, and materials science Frenkel & Smit (2001).While the atomic forces needed for MD simulation can be obtained on-the-fly via ab-initio quantum-mechanical (QM) calculations Car & Parrinello (1985), this is prohibitively expensive for realistic length and time scales Friesner (2005). Machine learning force fields (MLFFs) have recently emerged as a promising option to serve as surrogate models for QM calculations, demonstrating the ability to capture complex many-body interactions, and in some cases transfer flexibly across chemical space Schütt et al. (2018); Hu et al. (2021); Liu et al. (2022); Gasteiger et al. (2021; 2020); Batzner et al. (2022); Musaelian et al. (2022); Batatia et al. (2022); Schütt et al. (2021). Graph neural network (GNN)-based MLFFs trained on large ab-initio datasets are increasingly being used to model challenging and important chemical systems with favorable results Merchant et al. (2023); Kozinsky et al. (2023); Chen & Ong (2022); Schaarschmidt et al. (2022); Takamoto et al. (2022); Majewski et al. (2023); Charron et al. (2023); Batatia et al. (2023); Kovács et al. (2023); Shoghi et al. (2023); Gasteiger et al. (2022); Deng et al. (2023).
@@ 1 Introduction
Figure 1: Machine learning force field (MLFF) failure modes. (a) Illustrative examples of true and learned potential energy surfaces (PES) and resulting dynamics for unstable MLFFs. MLFFs can be accurate in approximating much of the PES, but contain regions where energy and forces estimates deviate significantly from the true PES, leading to sampling of highly unphysical regimes. As a result, observables computed from MD simulation may be biased by the oversampling of unphysical states, or have high statistical error in the extreme case of unrecoverable simulation collapse. (b) Examples of stable versus unstable configurations sampled by MLFFs during molecular dynamics simulation of systems considered in this work. (c) Selected states from an unstable MD trajectory of aspirin.
MD simulations aim to accurately estimate system observables like the radial distribution function, virial stress tensor, and diffusivity coefficient. This often requires long simulation to fully explore the underlying PES . Unfortunately, MLFFs are known to produce unstable simulations, meaning that they can irreversibly enter unphysical regions of phase space (e.g., a bond breaking event in a non-reactive system at low temperature) Fu et al. (2022); Stocker et al. (2022); Vita & Schwalbe-Koda (2023); Bihani et al. (2023); Morrow et al. (2023); Wang et al. (2023b). Sampling of such regions can lead to inaccuracies in computed observables as the MLFF gradually drifts from the distribution of its training data. In extreme cases, instabilities can lead to unrecoverable simulation collapse, in which case computed observables may have high statistical error due to insufficient sampling. In either case, instability can limit the ability of MLFF-based MD simulations to investigate long-timescale phenomena like ion diffusion and protein folding, as well as rare events that may require extensive sampling to encounter. Figure 1 illustrates typical MLFF instability behaviors and provides examples of unstable structures sampled during MD simulation.
Simulation instability has been shown to have an unreliable correlation with the energy and force error metrics typically used to train and evaluate MLFFs Fu et al. (2022); Stocker et al. (2022); Bihani et al. (2023). Recent works have introduced alternative simulation-based objectives, such as localized Wang et al. (2023b) or reweighted Ge et al. (2024) energy and force errors, to achieve better downstream stability. However, these approaches do not allow the MLFF to visit new configurations, limiting the distribution over which it is trained and thus the potential improvement. Alternatively, expanding the phase space coverage of the dataset can be an effective way to combat MLFF instability Fu et al. (2022); Stocker et al. (2022); Vita & Schwalbe-Koda (2023); Bihani et al. (2023); Morrow et al. (2023); Wang et al. (2023b). This is typically accomplished via active learning Smith et al. (2018); Vandermause et al. (2020); Schran et al. (2020); Lin et al. (2021); Kulichenko et al. (2023) approaches, where new atomistic configurations are selected, ab-initio QM calculations are performed, and the MLFF is retrained on the expanded dataset. Howver, these techniques rely on performing additional ab-initio calculations to expand the dataset. With MLFFs being trained on increasingly diverse datasets Shoghi et al. (2023); Batatia et al. (2023); Kovács et al. (2023) and used to simulate larger atomistic systems, the expense of these calculations may hamper the practicality of active learning workflows. This suggests the need for new training strategies or additional sources of information beyond energies and forces to train stable MLFFs.
@@
In this work, we bridge this gap by using both system observables and ab-initio QM data to improve MLFF stability. We introduce Stability-Aware Boltzmann Estimator (StABlE) Training, which uses efficient, parallelized MD simulations to rapidly explore regions of molecular phase space where the MLFF becomes unstable, followed by a targeted refinement of these regions using reference system observables. The key to efficient and numerically stable training lies in the Boltzmann Estimator, which enables end-to-end gradient-based learning without backpropagating directly through MD simulations. We also introduce a localized version of the Boltzmann Estimator, which enables targeted refinement of local instabilities. This is particularly important for stabilizing simulations of large, condensed-phase systems. StABlE Training is a self-contained, efficient process that leverages learning signals from both reference observables and existing QM data, with no reliance on performing additional QM reference calculations.
@@
We demonstrate StABlE Training on three systems and MLFF architectures: 1) simulation of aspirin with SchNet Schütt et al. (2018), 2) simulation of the Ac-Ala3-NHMe tetrapeptide with NequIP Batzner et al. (2022), and 3) simulation of an all-atom water system with GemNet-T Gasteiger et al. (2021). Relative to MLFFs trained solely on energies and forces, our StABlE-trained models produce significantly more stable MD simulations, recover observables more accurately, exhibit better generalization to unseen simulation temperatures, and outperform models trained on 50 times larger, labeled datasets. Our results suggest that utilizing both quantum-mechanical and observable-based modalities is required to fully exploit the available learning signal in reference datasets and train stable and accurate MLFFs. To our knowledge, StABlE Training is the first method that combines these data modalities to improve the stability of neural network potentials in MD simulations.
@@ 2 Preliminaries
Molecular dynamics simulation is used to evolve the positions and momenta of an atomistic system. Given a system of $N$ atoms, its state at time $t$ is defined by $\Gamma(t)=\{r(t),p(t)\}$, where $r(t),p(t)\in\mathbb{R}^{N\times 3}$ are the position and momenta of the atoms. We assume the system has a scalar-valued Hamiltonian function $\mathcal{H}:\mathbb{R}^{N\times 3}\times\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}$ of the form $\mathcal{H}(\Gamma)=\sum_{i=1}^{N}\frac{p^{{(i)}^{2}}}{2m^{(i)}}+U(r)\,$ where $U:\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}$ is a potential energy function and $m^{(i)}$ and $p^{(i)}$ are the mass and momentum of atom $i$. By updating in the direction of the per-atom forces $-\frac{\partial U}{\partial r}$ using a numerical integration scheme such as Langevin dynamics Bussi & Parrinello (2007), a sequence of $K$ simulation states $\{\Gamma(t)\}_{t=0}^{K}$ is produced.
Machine Learning Force Fields.
A MLFF is a function approximator which learns a potential energy $U_{\theta}$ and forces $F_{\theta}=-\nabla_{r}U_{\theta}=(F_{\theta}^{(1)},\dots,F_{\theta}^{(n)})$ from QM reference data. MLFFs are trained to minimize the following regression loss, with supervision from a dataset of reference energies and forces $\mathcal{D}_{train}=\{(r_{i},U_{i},F_{i})\}_{i=1}^{N}$.
@@
$\mathcal{L}_{QM}=\frac{1}{N}\sum_{i=1}^{N}\left[\lambda_{U}|U_{i}(\Gamma)-U_{\theta}(\Gamma)|^{2}+\lambda_{F}\sum_{j=1}^{n}\|F_{i}^{(j)}(\Gamma)+\nabla_{r^{(j)}}U_{\theta}(\Gamma)\|_{2}^{2}\ \right]$ (1)
System Observables.
@@
Observables, $g(\Gamma(t))$, characterize the state of a MD simulation at time $t$, and relate to macroscopic properties or experimental measurements of the system. Examples include the radial and angular distribution functions, velocity autocorrelation function, and diffusivity coefficient. Observables can be computed conveniently as an empirical average over states from a MD simulation. This is justified by the ergodicity hypothesis, under which a time-average over a sufficiently long simulation converges to a distributional average over the Boltzmann distribution. More details on the observables used in this work can be found in Supplementary Section A.4.
Training MLFFs with Observables.
Observables have been used extensively in the historical development of potentials Cornell et al. (1995); Marrink et al. (2007); Li et al. (2018), and more recently for MLFFs Thaler & Zavadlav (2021); Wang et al. (2020) in the context of condensed-phase Wang et al. (2023a) and titanium Röcken & Zavadlav (2023) systems, enhanced sampling of rare events Šípka et al. (2023), and protein folding simulation Ingraham et al. (2019); Navarro et al. (2023); Kolloff & Olsson (2023). Training with observables requires an efficient way to compute gradients through MD simulations while avoiding numerical instability and memory limitations. The approach is appealing due to the lack of reliance on expensive ab-initio quantum mechanical calculations, and the possibility of improved empirical consistency in settings where the underlying ab-initio method may be unreliable Cheetham & Seshadri (2024). However, past works do not consider MLFFs with stability problems, and dispense entirely with using QM data, thus losing the first-principles guarantees of matching an ab-initio PES.
Figure 2: Schematic of Stability-Aware Boltzmann Estimator (StABlE) Training procedure. Our proposed StABlE Training procedure begins with conventional pre-training on a small reference dataset of QM calculations. This dataset remains fixed throughout the procedure, and is never expanded with new calculations. Upon convergence of pre-training, StABlE alternates between two main phases, simulation and learning. In the simulation phase, we perform many molecular dynamics simulations in parallel with the MLFF and find regions of instability. When a sufficient fraction of simulations become unstable, we enter the learning phase, where the MLFF is further trained to match known system observables from ab-initio simulation or experiment. Gradients are computed efficiently through the MD simulation via our Boltzmann Estimator. After a sufficient reduction in the portion of unstable trajectories, we re-enter the simulation phase, and repeat the training cycle until a predetermined computational budget is reached.
@@ 3 Methods
We present StABlE Training, our proposed procedure to train stable and accurate MLFFs by leveraging both reference quantum mechanical data and system observables. Our approach is enabled by Boltzmann Estimators, which allows gradient-based optimization of MLFFs based on system observables.
@@ 3.1 Boltzmann Estimator
Reference system observables can be estimated from high-fidelity MD simulations or experimental measurements. Since observables are linked to the MLFF parameters through a molecular simulation, training with this source of information requires a reliable way to optimize through MD trajectories.
Formally, we define $g(\Gamma)$ to be a vector-valued observable of a state $\Gamma$, and $g_{\text{ref}}$ to be the reference value of the observable. To train our MLFF to match a reference observable, we minimize the following loss function:
@@
where $P_{\theta}(\Gamma)$ is the equilibrium distribution induced by the MLFF $U_{\theta}$. In this work, we primarily consider systems with a fixed volume, temperature, and number of particles, corresponding to the canonical (NVT) ensemble. We note, however, that the estimator is readily applicable to other ensembles, including the isothermal/isobaric (NPT) and grand canonical ($\mu VT$) ensembles (see Supplementary Sections A.2 and A.10 for details). In the canonical ensemble, microstates obey a Boltzmann distribution, $P_{\theta}(\Gamma)={\exp\left(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma)\right)}/{C(\theta)}$, where $\mathcal{H}_{\theta}(\Gamma)$ is the Hamiltonian, $T$ is the sampling temperature, $k_{B}$ is Boltzmann’s constant, and $C(\theta)=\int\exp(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma^{\prime}))\mathop{}\!\mathrm{d}\Gamma^{\prime}$ is the normalizing partition function. MD simulation is required at each training iteration to sample from this distribution and estimate the loss.
Optimizing the MLFF requires computing gradients of Equation 2 with respect to $\theta$. The required gradient, $\nabla_{\theta}\mathcal{L}_{\text{obs}}$, can be decomposed via the chain rule as follows:
$\nabla_{\theta}\mathcal{L}_{\text{obs}}^{\top}$ $=\frac{\partial\mathcal{L}_{\text{obs}}}{\partial{\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)[g(\Gamma)]}}}\frac{\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}{\partial\theta}$ $=2(\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]-g_{\text{ref}})^{\top}\frac{\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}{\partial\theta}\,.$ (3)
The non-trivial quantity to compute is the Jacobian, ${\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}/{\partial\theta}$. One way to estimate it is by using a chain rule expansion that corresponds to each step of the unrolled MD simulation Wang et al. (2020). The adjoint method Chen et al. (2019) can be used to limit the memory footprint at the expense of increased computation, but can still lead to numerical instability for long trajectories Wang et al. (2023a); Šípka et al. (2023).
We can avoid direct backpropagation through the simulation by noting that the equilibrium state distribution $P_{\theta}(\Gamma)$ is independent of the algorithm (i.e., MD integrator) used to sample the distribution. This is analogous to implicit differentiation techniques for differentiable optimization Amos & Kolter (2017); Gould et al. (2016); Ren et al. (2022); Blondel et al. (2022); Négiar et al. (2023), in which the solution to an optimization problem is decoupled from the numerical solver used to obtain it. We leverage the known Boltzmann form of $P_{\theta}(\Gamma)$ to construct an unbiased estimator of the Jacobian.
@@
Given $N$ independent samples $\Gamma_{1},\ldots,\Gamma_{N}$ from a Boltzmann distribution $P_{\theta}(\Gamma)$, we define the $N$-sample Boltzmann estimator $\mathcal{E}(\Gamma_{1},\ldots,\Gamma_{N})$ of the Jacobian ${\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}/{\partial\theta}$ as, $\mathcal{E}(\Gamma_{1},\ldots,\Gamma_{N})=$ $\frac{N}{k_{B}T(N-1)}\left[\hat{\mathbb{E}}\left[g(\Gamma)\right]\hat{\mathbb{E}}\big{[}{\nabla_{\theta}U_{\theta}(\Gamma)}\right]^{\top}\!-\hat{\mathbb{E}}\left[{g(\Gamma)\cdot\nabla_{\theta}U_{\theta}(\Gamma)}^{\top}\right]\big{]}\,,$ (4) where $\hat{\mathbb{E}}[f(\Gamma)]=\frac{1}{N}\sum_{i=1}^{N}f(\Gamma_{i})$ denotes an empirical mean over the samples.
This estimator provides an unbiased estimate of the Jacobian $\frac{\partial\mathbb{E}_{\Gamma\sim P_{\theta}(\Gamma)}[g(\Gamma)]}{\partial\theta}$. A proof is provided in Supplementary Section A.1. The estimator is related to the REINFORCE trick Williams (1992) and policy gradient estimators from reinforcement learning Silver et al. (2014) when a Boltzmann state distribution is assumed. The result can also be derived using thermodynamic perturbation theory Zwanzig (1954); Thaler & Zavadlav (2021).